Selective Oxidation The standard reduction potential for the half-reaction Sn4+ + 2e - Sn2+ is +0.15 V. Consider data from thmedia%2F909%2F9095418a-1167-4704-8c55-f6
Answer #1

The standard reduction potential of the half reactions are -

Sn + + 2e → Sn2+ E° = 0.15V

Sn2+ + 2e + Sn E° = -0.14V

We know that when two half reactions are coupled, the half reaction with the higher reduction potential will undergo reduction, and that with the lower reduction potential will undergo oxidation.

Reduction will occur at cathode (right hand electrode) and oxidation will occur at anode (left hand electrode).

Ecell = Ecathode – Eanode

Condition for reaction - Reaction will only occur when sathode > EO , i.e. cell > 0

[According to modern conventions, we are considering standard reduction potential only. Oxidation potentials are not considered nowadays.]

Now according to the question, the required half reactions are cathodic, that means it will occur at cathode. So its reduction potential must be greater than that at the anode.

According to the question, there will be spontaneous oxidation of Sn to Sn2+.So the condition of redox reaction is fulfilled. So the cathodic half reaction will have Eo > -0.14 V.

Also, it is mentioned that there will be spontaneous oxidation of Sn2+ to Sn4+. So the condition of redox reaction is not fulfilled. So the cathodic half reaction will have Eo < 0.15 V.

So we have to find reactions whose standard reduction potentials are less than 0.15 V but more than -0.14 V.

-0.14 V < Eo < 0.15 V. Only then Eocell will be positive and the reaction will occur.

Given:

2H+ + 2e + H EⓇ = 0.00 V

Fe3+ + 3e + Fe Eº = -0.036 V

Sn2+ + 2e + Sn E = -0.14 V

Fe2+ + 2e + Fe E° = -0.44 V

Pb2+ + 2e + Pb Eº = -0.13 V

Ni2+ + 2e + Ni Eº = -0.23 V

The half cell reactions that are satisfying the criteria -0.14 V < Eo < 0.15 V are -

1) 2H+ + 2e + H EⓇ = 0.00 V

2)Fe3+ + 3e + Fe Eº = -0.036 V

3) Pb2+ + 2e + Pb Eº = -0.13 V

These 3 are the required answer.

(The reaction Sn2+ + 2e + Sn E = -0.14 V will not be considered as it is not more than -0.14 V).